OpenStudy (anonymous):

how about this one.. v=1/3Bh, B=28.00, h=7 I am lost.

OpenStudy (lgbasallote):

where are you getting confused?

OpenStudy (anonymous):

just not sure how to solve it

OpenStudy (lgbasallote):

okay it says \[V = \frac 13 Bh\] it says B = 28.00 h = 7 so just put that in \[V = \frac 13 \times 28 \times 7\]

OpenStudy (lgbasallote):

do you get it @cacotto ?

OpenStudy (anonymous):

I skipped it, I will come back to it

OpenStudy (lgbasallote):

okay